# LLM-as-a-qualitative-judge: automating error analysis in natural language generation

### v2

@@ Abstract
Prompting large language models (LLMs) to evaluate generated text, known as LLM-as-a-judge, has become a standard evaluation approach in natural language generation (NLG), but is primarily used as a quantitative tool, i.e. with numerical scores as main outputs. In this work, we propose LLM-as-a-qualitative-judge, an LLM-based evaluation approach with the main output being a structured report of common issue types in the NLG system outputs. Our approach is targeted at providing developers with meaningful insights on what improvements can be done to a given NLG system and consists of two main steps, namely open-ended per-instance issue analysis and clustering of the discovered issues using an intuitive cumulative algorithm. We also introduce a strategy for evaluating the proposed approach, coupled with $\sim$300 annotations of issues in instances from 12 NLG datasets. Our results show that LLM-as-a-qualitative-judge correctly recognizes instance-specific issues in 2/3 cases and is capable of producing error type reports resembling the reports composed by human annotators. Our code and data are publicly available [^1] Code & data: https://github.com/tunde-ajayi/llm-as-a-qualitative-judge.
@@ 1 Introduction
Prompting large language models (LLMs) to output evaluation scores, known as LLM-as-a-judge (Zheng et al., 2023a), has become a standard approach for evaluating performance in natural language generation (NLG) tasks. In contrast to classic statistical measures such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), or METEOR (Banerjee and Lavie, 2005), which primarily rely on surface-level lexical overlap, LLM-as-a-judge evaluates text based on semantic understanding, allowing it to better handle diverse phrasings that convey equivalent meanings. Consequently, it shows stronger alignment with human judgment in various tasks, including machine translation (Kocmi and Federmann, 2023), summarization (Clark et al., 2023), or open-ended instruction following (Ye et al., 2024), especially with strong recent LLMs.
@@
Figure 2: Illustration of the proposed approach and evaluation methodology.
Recent works propose various extensions of the LLM-as-a-judge approach, including pairwise model comparison (Zheng et al., 2023b), finetuning LLMs for evaluation (Kim et al., 2024b, c), multi-criteria evaluation (Liang et al., 2023; Fu et al., 2024), the dynamic selection of evaluation criteria (Ye et al., 2024), or even using per-instance evaluation checklists (Cook et al., 2024; Kim et al., 2025). A common technique to improve LLM-as-a-judge is to ask a model to output an explanation for the predicted score(s).
However, even with the extensions listed above, LLM-as-a-judge remains primarily a quantitative evaluation tool, i.e., the final result used by researchers and practitioners is quantitative evaluation scores. At the same time, language generation is a complex multi-faceted task with a vast space of potential issues, including in various aspects of generated texts (grammaticality, factuality, logical coherence, etc), in preprocessing of the input data and postprocessing of the NLG outputs, or even with user requests. An effective and commonly used strategy for spotting such issues is a manual qualitative error analysis of a subset of predictions, which allows developers to identify artifacts, fix system issues, and detect flaws in quantitative evaluation. Yet this analysis is often skipped in practice, due to overreliance of developers on quantitative metrics, as well as high demand in terms of time and effort needed to conduct such analysis.
In this work, we introduce LLM-as-a-qualitative-judge, a novel approach which automates error analysis, with the main output being a a structured report aggregating the common qualitative error types in the NLG outputs for a given dataset. The two key steps of LLM-as-a-qualitative-judge are (1) open-ended per-instance error analysis and (2) clustering of the discovered error types. Per-instance analysis implies prompting an LLM to detect an issue in the given NLG system output, where an issue may be arbitrary, i.e. we do not provide any predefined set of possible issues. For error clustering, we propose an intuitive and effective algorithm which resembles how humans solve the corresponding task. The examples of the generated reports are presented in Figure 1, and a high-level illustration of the proposed approach is presented in Figure 2.
To evaluate the proposed LLM-as-a-qualitative-judge, we manually annotate error types in a diverse set of 12 generative datasets (7 tasks and 6 domains for one of the tasks) and introduce an LLM-as-a-judge-based strategy for comparing the errors detected by an LLM and by a human expert. We also manually annotate potential clustering of errors in each dataset, which allows us to evaluate the automatically produced clustering, by calculating its agreement with the human-annotated clustering.
@@
  • We collect a set of $\sim$300 manual annotations of open-ended issues in the instances coming from 12 diverse NLG datasets, as well as the manual annotations of their per-dataset clustering;
  • We introduce a strategy for evaluating LLM-as-a-qualitative-judge and show that LLM-as-a-qualitative-judge is capable of producing error type reports which resemble the reports produced by humans.
We hope that the proposed LLM-as-a-qualitative-judge approach will reduce the time and effort required for issue analysis and will help practitioners to more easily spot issues in their NLG pipelines.
@@ 2 Proposed approach
For the purposes of our algorithm, the dataset consists of instances, each represented by a user input (a string containing a task instruction and the input data), a ground truth response (a string defining a correct answer), and a generated response (a final output of the NLG system). Each instance can be optionally augmented with other fields, e.g., the intermediate outputs of an NLG system such as retrieved documents in retrieval-augmented generation, or additional information on the NLG system, e.g., a definition of a task metric. The LLM-as-a-qualitative-judge algorithm is summarized in Algorithm 1, illustrated in Figure 2 and described step-by-step below.
@@
For each instance, we prompt an LLM to identify “one, most important, specific, clearly visible issue”, provided with a user input, a ground truth response, a generated response, and optionally other fields as described above. We prompt an LLM to output a detailed analysis of a given instance, followed by a special separator and a final 1–2 sentence description of an identified issue, which is referred to as a per-instance issue explanation in the following steps of the approach. The particular prompt used for per-instance analysis is presented in App. Figure 5.
@@
1:a list of user inputs $U$, a list of ground truth responses $R^{\mathrm{gt}}$, a list of generated responses $R$ — all of length $N$;
@@
[…]
25:return $C$, $A$
@@
The second step in LLM-as-a-qualitative-judge consists of clustering issues discovered in the first step and forming a final report of main issue types based on the clustering results. This can be, in principle, done with any clustering algorithm, e.g., k-means with BERT-based embeddings (Devlin et al., 2019) or directly prompting a strong LLM to output clustering (Viswanathan et al., 2024), provided with clustering inputs in a single prompt. In the experiments, we demonstrate the downsides of these approaches, e.g., classic approaches perform poorly on our data, and clustering with direct prompting fails for larger datasets, weaker LLMs, and does not ensure the structural correctness of the generated report.
Inspired by how humans would cluster issues, we propose an intuitive cumulative issue clustering algorithm. Our clustering algorithm goes through instances one-by-one and gradually builds the issue types report. For each instance, we provide an LLM with the current report and the per-instance issue explanation, and prompt the LLM to decide whether this issue explanation can be attributed to one of the already discovered issue types (clusters) or it should form a new issue type (cluster). In the former case, we augment the counter of the corresponding issue type by one. In the latter case, we also prompt an LLM to formulate a short name and a 1–2 sentence description of a new error type, based on the per-instance issue description. In particular, we instruct an LLM to formulate “a fine-grained issue type that can be generalized to other instances”. The new issue type is then added to the report, represented by the generated issue type name, description, and the counter set to one. The first instance in the dataset is always a new issue type. Appendix Figures 7 and 8 show the prompts used for the two described clustering steps.
@@
3 Evaluation protocol
In this section, we introduce an evaluation protocol to assess the effectiveness of both steps in the LLM-as-a-qualitative-judge approach.
Per-instance analysis.
The core of our evaluation strategy is to collect manual annotations of failure cases for a set of instances from various tasks and domains. For each instance, consisting of a user input, a ground truth response, a generated response, a task metric, and optionally retrieved documents, an annotator’s task is to formulate what the particular issue is in this instance. We then prompt an evaluator LLM to judge whether the issue explanation determined by the LLM-as-a-qualitative-judge for a particular instance matches the issue determined by the human annotator. The outputs from the evaluator LLM are binary and can be accumulated into a per-instance analysis accuracy score.
@@ 3 Evaluation protocol
Instruction following FLASK Ye et al. (2024) 34
Translation en-ru WMT’22 Kocmi et al. (2022) 38
Long context QA Elitr-Bench Thonet et al. (2025) 26
Semantic parsing PIZZA Arkoudas et al. (2022) 34
Grade school math GSM8K Cobbe et al. (2021) 17
Detoxification ParaDetox Dementieva et al. (2024) 36
@@
Factoid QA in Russian MKQA (ru) Longpre et al. (2021) 29
Biomedical QA BioASQ Krithara et al. (2023) 27
Lifestyle forum QA RobustQA Han et al. (2024); Santhanam et al. (2022) 21
Writing forum QA 13
Search engine queries SearchQA Dunn et al. (2017) 13
Educational QA SyllabusQA (Fernandez et al., 2024) 9
@@
Dataset composition.
Table 1 provides the data summary. We consider a diverse pool of 7 generative tasks, and for one of the tasks, namely retrieval-augmented question answering (RA-QA), we consider a diverse pool of 6 domains. For each considered dataset, we manually label failures in up to 40 generations from one of the open-source LLMs (Qwen-2.5-1b, Llama-3.2-1b, Command-R-35b, Vicuna-1.5-13B). All the labeling was performed by the authors of the paper in Google Spreadsheets [^3] https://docs.google.com/spreadsheets. The final dataset comprises 297 instances.
4 Experiments
@@ 4.1 Experimental setup
We test per-instance analysis with a range of commercial and open-source LLMs. For issue clustering, we compare the proposed cumulative clustering approach to the direct LLM prompting and classic clustering approaches. All clustering runs operate on the per-instance analysis output by GPT-4o, and clustering results are averaged over 3 runs from different random seeds. We test issue clustering with three LLMs: GPT-4o, Gemini-2-Flash, and Qwen-2.5-7B. Tables 5 and 6 in Appendix list references and licenses of the used models and datasets, respectively.
Figure 3: Examples of per-instance analysis.
Open-source LLMs are run on a single V100 GPU with greedy decoding ($\sim$20 GPU-hours in total). Commercial LLMs are run via API with requesting json output format. Prompts used for all the stages are presented in the Appendix A. Exact task formulations in prompts were adjusted by only using three RA-QA datasets (MKQA (ru), RobustQA Lifestyle and Writing).
For each instance, LLM-as-a-qualitative-judge is provided with a user input, a ground truth response, a generated response, 5 retrieved documents (only for RA-QA), and a short task comment. The task comment describes the task metric (in one sentence), provides a comment on the nature of ground truth responses (either that it is the expected answer or that it is only one of the possible correct answers), and also contains a comment that retrieval-augmented generation (RAG) or Chain-of-Thought (COT) prompting was used, when applicable (6 datasets with RAG and 2 datasets with COT). The used task metrics are binary LLM-as-a-judge (the generated response is accepted or not) or binary Match (outputs True is one of the ground truth answers is contained a substring in the generated response, and False otherwise). Task comments for all datasets are also presented in Appendix E.
For classic clustering approaches, we use the scikit-learn implementation with BERT embeddings and tune hyperparameters as described in App. B. For methods requiring the number of clusters, we set it the same as in the annotator’s data.
Meta-evaluation of an evaluator LLM.
To ensure the reliability of the evaluator LLM, we collected a small meta-evaluation dataset of 50 instances from 4 datasets (MKQA (ru), RobustQA Writing, FLASK, MultiDetox), where the equivalence of the LLM-as-a-qualitative-judge’s and human annotator’s per-instance analysis was judged by a human annotator and can be compared to the evaluator LLM’s verdict. Strong commercial LLMs, such as GPT-4o, Gemini-2.0-Flash, and Claude-3.7-Sonnet, achieved a meta-evaluation accuracy of 85-90% on this dataset, and an open-source Solar-10.7B (Kim et al., 2024a) achieved a meta-accuracy of 60%. In all the experiments, we use claude-3-7-sonnet-20250219 as the evaluator LLM, to avoid using the same LLM for analysis and for evaluation.
@@
4.2 Per-instance analysis
@@
Figure 3 shows examples of issue explanations generated by GPT-4o and Qwen-2.5-7B. In the manual inspection of the generated issue explanations, we observed correct explanations for various kinds of issues, and rows 1–3 demonstrate such examples. We also highlighted three common groups of mistakes. The first repeated problem in per-instance analysis is logical issues. In the example in row 4, the issue is that the ground truth response is not contained as a verbatim substring in the model-generated response, which is a definition of a task metric. However, both GPT-4o and Qwen-2.5-7B claim that the failure in the substring match is caused by the model response containing extra generated information. Such an explanation logically contradicts the task metric, i.e., extra content can only increase chances of finding a given substring in the response, but cannot be a reason for its absence.
The second common problem in per-instance analysis is the oversimplification of an issue, especially for more unexpected issues, such as an ambiguous user input or an error in evaluation. In the example in row 5, the issue is an ambiguous user question, i.e., both the ground truth and the generated response are correct and provide two different interpretations of the user question. However, GPT-4o and Qwen-2.5-7B report the over-simplified issue of the generated response not providing the same answer as the ground truth response.
Finally, the third common reason for a per-instance issue explanation not being accepted by the evaluator LLM, is the subjectivity of some issues in the dataset. For example, a human-annotated issue in row 6, “The generation was stopped too early because of the reached maximum new tokens limit”, is evaluated to be not equivalent to the LLM-as-a-qualitative-judge-generated issue “The generated response provides an incomplete overview <…>”. While these two issue explanations are indeed different, they are both correct, and the LLM-generated explanation follows from the annotator’s explanation.
Future work could attempt to alleviate the first two mentioned problems by using stronger reasoning approaches in per-instance analysis, and the third problem by improving an evaluator LLM.

### v4

@@ Abstract
Prompting large language models (LLMs) to evaluate generated text, known as LLM-as-a-judge, has become a standard evaluation approach in natural language generation (NLG), but is primarily used as a quantitative tool, i.e. with numerical scores as main outputs. In this work, we propose LLM-as-a-qualitative-judge, an LLM-based evaluation approach with the main output being a structured report of common issue types in the NLG system outputs. Our approach is targeted at providing developers with meaningful insights on what improvements can be done to a given NLG system and consists of two main steps, namely open-ended per-instance issue analysis and clustering of the discovered issues using an intuitive cumulative algorithm. We also introduce a strategy for evaluating the proposed approach, coupled with $\sim$300 annotations of issues in instances from 12 NLG datasets. Our results show that instance-specific issues output by LLM-as-a-qualitative-judge match those annotated by humans in 2/3 cases, and that LLM-as-a-qualitative-judge is capable of producing error type reports resembling the reports composed by human annotators. We also demonstrate in a case study how the use of LLM-as-a-qualitative-judge can substantially improve NLG systems performance. Our code and data are publicly available [^1] Code & data: https://github.com/tunde-ajayi/llm-as-a-qualitative-judge.
@@ 1 Introduction
Prompting large language models (LLMs) to output evaluation scores, known as LLM-as-a-judge (zheng23-judgingllmasajudgemtbench), has become a standard approach for evaluating performance in natural language generation (NLG) tasks. In contrast to classic statistical measures such as BLEU (papineni-etal-2002-bleu), ROUGE (lin-2004-rouge), or METEOR (banerjee-lavie-2005-meteor), which primarily rely on surface-level lexical overlap, LLM-as-a-judge evaluates text based on deep semantic understanding, allowing it to better handle diverse phrasings that convey equivalent meanings. Consequently, it shows stronger alignment with human judgment in various tasks, including machine translation (kocmi-federmann-2023-large), summarization (seahorse), or open-ended instruction following (flask), especially with strong recent LLMs.
@@
Figure 2: Illustration of the proposed LLM-as-a-qualitative-judge approach.
Recent works propose various extensions of the LLM-as-a-judge approach, including pairwise model comparison (zheng2023judging), finetuning LLMs for evaluation (prometheus; prometheus2), multi-criteria evaluation (liang2023holistic; fu-etal-2024-gptscore), the dynamic selection of evaluation criteria (flask), or even using per-instance evaluation checklists (checklists; biggen). A common technique to improve LLM-as-a-judge is to ask a model to output an explanation for the predicted score(s).
However, even with the extensions listed above, LLM-as-a-judge remains primarily a quantitative evaluation tool, i.e., the final result used by researchers and practitioners is quantitative evaluation scores. At the same time, language generation is a complex multi-faceted task with a vast space of potential issues, including in various aspects of generated texts (grammaticality, factuality, logical coherence, etc.), in preprocessing of the input data and postprocessing of the NLG outputs, or even with user requests. An effective and commonly used strategy for spotting such issues is a manual qualitative error analysis of a subset of predictions, which allows developers to identify artifacts, fix system issues, and detect flaws in quantitative evaluation. Yet this analysis is often skipped in practice van-miltenburg-etal-2023-barriers; van-miltenburg-etal-2021-underreporting, due to overreliance of developers on quantitative metrics, as well as high demand in terms of time and effort needed to conduct such analysis.
In this work, we introduce LLM-as-a-qualitative-judge, a novel approach which automates error analysis, with the main output being a a structured report aggregating the common qualitative error types in the NLG outputs for a given dataset. The two key steps of LLM-as-a-qualitative-judge are (1) open-ended per-instance error analysis and (2) clustering of the discovered error types. Per-instance analysis implies prompting an LLM to detect an issue in the given NLG system output, where an issue may be arbitrary, i.e. we do not provide any predefined set of possible issues. For error clustering, we propose an intuitive and effective algorithm which resembles how humans solve the corresponding task. An example of the generated report is presented in Figure 2 (left), and a high-level illustration of the proposed approach is presented in Figure 2 (right).
@@
  • In a case study on BigBenchHard tasks, we demonstrate that LLM-as-a-qualitative-judge can substantially improve the performance of NLG systems;
  • We collect $\sim$300 manual annotations of open-ended issues in the instances coming from 12 diverse NLG datasets, as well as the manual annotations of their per-dataset clustering;
  • We introduce a strategy for meta-evaluating LLM-as-a-qualitative-judge and show that LLM-as-a-qualitative-judge is capable of producing error type reports which resemble the reports produced by humans.
We hope that the proposed LLM-as-a-qualitative-judge approach will reduce the time and effort required for issue analysis and will help practitioners to more easily improve their NLG pipelines. Our code and data are available as https://github.com/tunde-ajayi/llm-as-a-qualitative-judge.
@@ 2 Proposed approach
For the purposes of our algorithm, the dataset consists of instances, each represented by a task input (a string containing a task instruction and the input data), a ground truth response (a string defining a correct answer), and a generated response (a final output of the NLG system). Each instance can be optionally augmented with other fields, e.g., the intermediate outputs of an NLG system such as retrieved documents in retrieval-augmented generation, or additional information on the NLG system, e.g., a definition of a task metric. The LLM-as-a-qualitative-judge algorithm is summarized in Algorithm 1, illustrated in Figure 2 and described step-by-step below.
@@
For each instance, we prompt an LLM to identify “one, most important, specific, clearly visible issue”, provided with a task input, a ground truth response, a generated response, and optionally other fields as described above. We prompt an LLM to output a detailed analysis of a given instance, followed by a special separator and a final 1–2 sentence description of an identified issue, which is referred to as a per-instance issue explanation in the following steps of the approach. The particular prompt used for per-instance analysis is presented in App. Figure 6.
@@
1:a list of task inputs $U$, a list of ground truth responses $R^{\mathrm{gt}}$, a list of generated responses $R$ — all of length NN;
@@
24:   $C[K][\mathrm{``count"}]\leftarrow C[K][\mathrm{``count"}]+1$
25:return $C$, AA
Figure 3: Case study on three BigBenchHard tasks: after building a simple pipeline for a task, we perform two rounds of generating issue reports with LLM-as-a-qualitative-judge (a table with issue types and their counts) and manually revising the pipeline based solely on the generated reports. Task performance is improved in all cases.
@@
The second step in LLM-as-a-qualitative-judge consists of clustering issues discovered in the first step and forming a final report of main issue types based on the clustering results. This can be, in principle, done with any clustering algorithm, e.g., k-means with BERT-based embeddings (bert) or directly prompting a strong LLM to output clustering (DBLP:journals/tacl/0002GGLWN24), provided with clustering inputs in a single prompt. In the experiments, we demonstrate the downsides of these approaches, e.g., classic approaches perform poorly on our data, and clustering with direct prompting fails for larger datasets, weaker LLMs, and does not ensure the structural correctness of the generated report.
Inspired by how humans would cluster issues, we propose an intuitive cumulative issue clustering algorithm. Our clustering algorithm goes through instances one-by-one and gradually builds the issue types report. For each instance, we provide an LLM with the current report and the per-instance issue explanation, and prompt the LLM to decide whether this issue explanation can be attributed to one of the already discovered issue types (clusters) or it should form a new issue type (cluster). In the former case, we augment the counter of the corresponding issue type by one. In the latter case, we also prompt an LLM to formulate a short name and a 1–2 sentence description of a new error type, based on the per-instance issue description. In particular, we instruct an LLM to formulate “a fine-grained issue type that can be generalized to other instances”. The new issue type is then added to the report, represented by the generated issue type name, description, and the counter set to one. The first instance in the dataset is always a new issue type. Appendix Figures 8 and 9 show the prompts used for the two described clustering steps.
@@
3 Case study
Our first set of experiments is targeted at demonstrating a practical utility of the proposed LLM-as-a-qualitative-judge.
Experimental setup.
We pick three tasks from a BigBenchHard collection (bbh), namely Date understanding, Word sorting, and Movie recommendation. For each dataset, we build a simplest pipeline consisting of prepending a simple system prompt “You are a helpful assistant. Output your answer after a final separator ‘Answer:’”, LLM generation with default hyperparameters, and a string match-based evaluation function. We then perform two rounds of generating an issue report with LLM-as-a-qualitative-judge (GPT-4.1) and manually revising the pipeline solely based on the generated report (issue types, their counts, and possibly 1 example of each issue type). More details are given in Appendix B.
Results.
As shown in Figure 3, the task performance is improved in all three cases. For example, in the Date understanding task, revisions inspired by the generated issue reports include explaining the date format in the prompt, suggesting to begin the response with determining a reference point in time, and providing a specific template for the output. These revisions improved performance from 4% to 62%.
4 Meta-evaluation methodology
This section described a methodology that we propose to meta-evaluate LLM-as-a-qualitative-judge. The corresponding set of experiments aims both to assess the effectiveness of two steps and to identify the optimal configurations for LLM-as-a-qualitative-judge.
Real-world data.
We manually annotate per-instance issues and their per-dataset clustering for a diverse pool of 12 datasets, with various open-source LLMs as generators. We consider 7 generative tasks, and for one of the tasks, namely retrieval-augmented question answering (RA-QA), we consider 6 domains. All the labeling was performed by the authors of the paper. The final dataset comprises 297 instances. Table 1 provides the data summary. More details on data annotation are presented in Appendix A.
Synthetic data.
Metrics.
For per-instance analysis, we prompt an evaluator LLM to judge whether the issue explanation determined by the LLM-as-a-qualitative-judge for a particular instance matches the issue determined by the human annotator. The outputs from the evaluator LLM are binary and are accumulated into a per-instance analysis accuracy score.
@@ 4 Meta-evaluation methodology
Instruction following FLASK flask 34
Translation en-ru WMT’22 kocmi-etal-2022-findings 38
Long context QA Elitr-Bench DBLP:conf/coling/ThonetBR25 26
Semantic parsing PIZZA DBLP:journals/corr/abs-2212-00265 34
Grade school math GSM8K DBLP:journals/corr/abs-2110-14168 17
Detoxification ParaDetox dementieva2024overview 36
@@
Factoid QA in Russian MKQA (ru) longpre-etal-2021-mkqa 29
Biomedical QA BioASQ krithara2023bioasq 27
Lifestyle forum QA RobustQA DBLP:conf/emnlp/Han00XWLWMC24; lotte 21
Search engine queries SearchQA DBLP:journals/corr/DunnSHGCC17 13
Educational QA SyllabusQA (DBLP:conf/acl/FernandezSL24) 9
@@
5 Meta-evaluation experiments
@@ 5.1 Experimental setup
We test per-instance analysis with a range of commercial and open-source LLMs, and issue clustering with three LLMs: GPT-4o, Gemini-2-Flash, and Qwen-2.5-7B. For issue clustering, we compare the proposed cumulative clustering approach to the direct LLM prompting and classic clustering approaches. All clustering runs operate on the per-instance analysis output by GPT-4o, and clustering results are averaged over 3 runs.
Tables 4 and 5 in Appendix list references and licenses of the used models and datasets, respectively. Prompts used for all the stages are presented in the Appendix G. Exact task formulations in prompts were adjusted by only using three RA-QA datasets (MKQA (ru), RobustQA Lifestyle and Writing).
For classic clustering approaches, we use the scikit-learn implementation with BERT embeddings and tune hyperparameters as described in App. C. For methods requiring the number of clusters, we set it the same as in the annotator’s data.
More details on experimental settings and as well as results on meta-evaluating the evaluator LLM are presented in Appendix B.
Appendix H presents the examples of generated error type reports, per-instance issue explanations, and confusion matrices for all considered datasets. We also provide an archive with experimental results in the project repository [^3] A zip archive with experimental results is available at https://github.com/tunde-ajayi/llm-as-a-qualitative-judge/tree/main/data.
5.2 Per-instance analysis
We note that our results are consistent with previously reported findings in the literature regarding the typical level of agreement between LLM-based evaluations and human judgments. For example, FLASK reports the highest correlation between model-based evaluation and human labelers, of 68% (Table 1 in flask), or METAL reports the highest agreement between the LLM evaluators and human scores of 59-82% for the first three criteria in (metal, Table 3, English).
To sum up, our results demonstrate the high effectiveness of strong LLMs in open-ended issue explanation for generative tasks. For practical applications, we recommend using recent models such as GPT-4o, Gemini-2.5-Flash or Qwen-2.5-32B.
@@ 5.2 Per-instance analysis
Figure 4: Examples of per-instance analysis.
@@
5.3 Examples of per-issue analysis
Figure 4 shows examples of issue explanations generated by GPT-4o and Qwen-2.5-7B. In the manual inspection of the generated issue explanations, we observed correct explanations for various kinds of issues, and rows 1–3 demonstrate such examples.
We also notice three groups of mistakes. The first occasional problem in per-instance analysis is logical issues. In the example in row 4, the issue is that the ground truth response is not contained as a verbatim substring in the model-generated response, which is a definition of a task metric. However, both GPT-4o and Qwen-2.5-7B claim that the failure in the substring match is caused by the model response containing extra generated information. Such an explanation logically contradicts the task metric, i.e., extra content can only increase chances of finding a given substring in the response, but cannot be a reason for its absence.